OpenStudy (floresf1):

Which function passes through the points (2, 15) and (3, 26)? y = 11x + 7 y = 11x − 7 y = 7x + 11 y = -11x − 7 y = 7x − 11

OpenStudy (anonymous):

use this formula\[y=mx+b\]where\[y=vertical~axis\]\[x=horizontal~axis\]\[m=slope\]\[b=y-intercept\] First find the slope using the given points Formula for slope is\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (floresf1):

ill try

OpenStudy (anonymous):

\[Let~say(2,15)=(x_1,y_1)\]\[Let~say(3,26)=(x_2,y_2)\]substitute this value into the slope formula

OpenStudy (anonymous):

\[m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

\[m=\frac{ 26-15 }{ 3-2 }\]

OpenStudy (anonymous):

\[m=11\]

OpenStudy (anonymous):

Now we have the value for x,y and m but we don't have b(y-intercept)

OpenStudy (anonymous):

I randomly choose \[(2,15)~as~(x,y)\]

OpenStudy (anonymous):

m=11

OpenStudy (anonymous):

Now substitute the value into this formula\[y=mx+b\]

OpenStudy (floresf1):

what does x equal

OpenStudy (anonymous):

\[15=11(2)+b\]\[15=22+b\]

OpenStudy (floresf1):

what does b equal to

OpenStudy (anonymous):

i randomly choose (2,15) to represent as(x,y)

OpenStudy (anonymous):

b=y-intercept

OpenStudy (floresf1):

so its what

OpenStudy (anonymous):

u have the value for x,y and m... Now substitute the value into\[y=mx+b\]to find the value of b

OpenStudy (anonymous):

The value for x is 2,y is 15 and m=11

OpenStudy (anonymous):

No we hv to find the value of b(y-intercept) to form the equation

OpenStudy (anonymous):

Now*

OpenStudy (anonymous):

\[15=2(11)+b\]\[15=22+b\]Subtract 22 on both sides

OpenStudy (anonymous):

\[15-22=22-22+b\]

OpenStudy (anonymous):

\[b=15-22\]\[b=?\]solve it @floresf1